# Policy Improvement using Language Feedback Models

### v3

@@ 4 Related Work
Instruction following in grounded environments has been explored in settings such as navigation (Chen & Mooney, 2011; Fried et al., 2018; Chen et al., 2019), game-playing (Andreas & Klein, 2015; Zhong et al., 2020), and robotics (Blukis et al., 2019; Shridhar et al., 2021a; Brohan et al., 2023). However, most prior work model environment observations separately from language instructions by using specialized encoders (e.g. ResNet (He et al., 2015)Bert (Devlin et al., 2019),  Clip (Radford et al., 2021)), then learn from data how to associate raw observations with language instructions. Instead of solely using raw observations, more recent work verbalize raw observations to describe environments in language (Shridhar et al., 2021b; Zhong et al., 2021; Schumann et al., 2024). In doing so, observations and instructions can be directly jointly reasoned over using language models to achieve more efficient and generalzable learning through large-scale pretraining. We build on this last direction by verbalizing raw observations into language descriptions to train language policies. However, unlike prior work that train language models to predict next actions, we develop language feedback models that critique verbalized observations and behaviour.
@@
LLMs exhibit an array of reasoning abilities by pretraining on vast quantities of text (Brown et al., 2020; Wei et al., 2022). A number of recent work investigate using LLMs as language agents to exploit this reasoning ability. Nakano et al. (2022)Yao et al. (2023) Deng et al. (2023) train instruction following language agents to interact with web browsers to answer questions or interact with web pages. Ahn et al. (2022) show that a language agent can be connected with verbalized robots via API interfaces for robotic control. While powerful, these prior work are limited in that they require querying an expensive LLM on-line. In contrast, our work examines settings where an LLM is not available on-line. Specially, we use LLMs to collect a small set of off-line data for training LFMs. The small and cost-effective LFMs are then used to identified desirable behaviour for on-line policy improvement without additional interactions with the LLM.
Table 1: Examples of verbalized environments. For brevity, we abbreviate long verbalized observations using “…”.
Benchmark
Context
Action
ALFWorld
Task: heat some egg and put it in diningtable. Observation: You arrive at loc 12. On the sinkbasin 1, you see… T-1 Observation: You are in the middle of a room… Action:go to sinkbasin 1 T-2 Observation: …
go to microwave 1
ScienceWorld
Task: Your task is to find a(n) living thing. First, focus on the thing. Then, move it to the purple box in the bathroom. Observation: You move to the kitchen. This room is called the kitchen. In it, you see: — the agent — a substance called air — a chair. On the chair is… In your inventory, you see: — an orange… T-1 Observation: The door is now open. Action: go to kitchen T-2 Observation… Action: open door to kitchen
open door to outside
Touchdown
Task: Follow the flow of traffic, with the row of flowers on your left and make a left at the intersection. There will be a white Billboard… Observation: behind you, you see: the right lane intersection, a large, blocky, gray… T-1 Observation: behind you, slightly… Action: slightly to your left …
straight ahead
@@
An important recent extension of language agents is to augment them with feedback. Ziegler et al. (2020)Stiennon et al. (2020), and Bai et al. (2022) learn reward models from human preference, which is then used to to learn a policy via reinforcement learning (RL). Instead of using human feedback, Bai et al. (2022) and Lee et al. (2023) use LLM feedback to train a separate reward model for RL for textual alignment. Huang et al. (2022) and Yao et al. (2023) use LLMs to reason about potential resolutions to failed actions. Yuan et al. (2024) use LLMs to generate new prompts and corresponding responses, then use an LLM reward model to identify good prompt-response pairs for self-improvement in text generation alignment. Unlike these approaches, we do not use LLMs during on-line policy improvement. We train an initial small language feedback model from offline LLM data, then use this small feedback model on-line during policy improvement. Additionally, we focus on-line improvement via language feedback for long-horizon, sparse reward, grounded environments instead of text generation alignment. Our procedure for batched, on-line imitation learning is similar to Dagger (Ross et al., 2011), which we compare to in Appendix C. However, we collect batched expert feedback to identify desirable behaviour instead of corrective actions.
@@ 5 Experiments and Analysis
We evaluate using Language Feedback Model for policy improvement on three distinct language grounding benchmarks. We compare this method against directly using LLMs as an expert policy for imitation learning. Formally, the environments from a benchmark are distinct partially-observed Markov Decision Processes that share some (or all) of the environment dynamics but have different instructions, observations, and/or action space.
@@ 5.1 Evaluation benchmarks
Table 1 shows examples of verbalized environments and tasks from each benchmark. Each benchmark provides distinct training and test environments to test generalization. In each environment, the agent takes actions to perform tasks outlined in a language instruction. The task is considered completed if and only if the agent solves the tasks within the preallocated number of steps. We evaluate using task-completion rate over test environments. The statistics from each benchmark is shown in Appendix A Table 6. These three benchmarks share challenges in sparse, delayed reward, partial observability, and compositional generalization to unseen tasks and environments.
ALFWorld is a verbalization of ALFRED (Shridhar et al., 2020), a natural language instruction following benchmark set in a 3D simulated kitchen. Here, the agent interacts with objects in kitchens to achieve compositional goals such as cleaning then microwaving potatoes. In ALFWorld (Shridhar et al., 2021b), raw state information from ALFRED are used to populate language templates that describe observations in language.
ScienceWorld is a textual simulation benchmark for basic science experiments (Wang et al., 2022). The agent interacts with objects to conduct experiments specified in natural language, such as determining the boiling temperature of a material. ScienceWorld is uniquely challenging to due the large amount of variations in task types (30), and parametric variations (10-1400) such as the specific substance to be melted. Furthermore, ScienceWorld has a substantially larger action space and longer horizon tasks.
Touchdown is a navigation benchmark where the agent navigates Google Street View images to follow long, compositional instructions (Chen et al., 2019). Touchdown requires jointly reasoning over natural images from Google Streetview with occlusion and multi-sentence natural language instructions that describe long-horizon goals. We introduce a new verbalization procedure for Touchdown based on matching noun-phrases and image patches with Clip embeddings to populate egocentric language templates. Behaviour cloning using our verbalization is detailed in Appendix A. Touchdown considers multiple subtasks, in this work we only test the agent’s ability to arrive at the correct location according to the instruction.
@@ 5.2 Methods
Table 2: LLM prompts used to collect desirable behaviour for imitation learning. ActPred uses LLMs to directly generate the appropriate action for each step, whereas Lfm uses LLMs to generate, in batch, feedback that identify which taken actions were productive. For brevity, we abbreviate long verbalized observations using “…”.
@@
Prompt
Your task is: look at alarmclock under the desklamp. You see: you are in the middle of a room. looking quickly around you, you see a bed 1, a desk 1, a drawer 17… what do you decide to do? available actions: examine shelf 1, examine shelf 2, go to bed… You decide to: go to desk 1. You see: you arrive at desk 1. what do you decide to do? available actions: examine desk 1… You decide to:
LLM Output
examine desk 1
@@
Prompt
You will be shown a playthrough for solving a task. Task: put two candle in drawer. Before: You open the drawer 6. The drawer 6 is open. In it, you see nothing. Step 21. Your action: close drawer 6. Result: You close the drawer 6… Step 22. Your action: put candle 3 in/on drawer 1. Result: You put the candle 3 in… Is the player on the right track to solve the task? Answer yes or no. If yes, list the helpful steps by the step number in bullet form.
LLM Output
Yes - Step 28 - Step 29…
We train BC baseline policies using existing demonstrations for each benchmark. We examine three different techniques for improving the BC policy. Table 2 shows examples of LLM prompts used for each technique.
@@
We learn a small and cost-effective feedback model described in Section 3.2 to identify desirable behaviour for imitation learning. First, we learn a feedback model on the training environments. Second, we use the feedback model to identify desirable behaviour in the training environments for policy improvement via imitation learning. To collect LLM feedback for training LFMs, we collect one rollout for each environment in a benchmark and sample 10k 20-step windows from the rollouts. Crucially, we limit the amount of feedback data collected from the LLM such that the number of output tokens produced by the LLM is identical to ActPred (we use 100k GPT-2 tokens for all benchmarks). This answers whether feedback model is a more cost-effective than direct action prediction for imitation learning.
LfmA: one-shot adaptation using feedback models.
@@ 5.3 Experiment details
We use the GPT-4 LLM (2023-03-15) for action prediction and feedback. We fine-tune the 770M Flan-T5 (Chung et al., 2022) to obtain policy and feedback models. We use descriptions of the most recent 20 steps as the verbalized observation $v$. All models are trained for 10k steps with batch size 20 and early stopping over validation demonstrations. Appendix E shows details on GPU usage.
@@
To train feedback models, we collect LLM feedback over 20-step windows. We then parse LLM feedback to identify whether the action taken in each step was productive to solving the tasks outlined in the instructions. We subsample the feedback data to obtain an even split of productive and not-productive actions. This data is split into a 80% train/20% validation dataset to train the LFM.
@@
To train policies, we fine-tune language models to minimize token-wise cross-entropy of the ground-truth verbalized action. During inference time, we consider a (potentially very large) set of plausible actions given by the environment. For each action, we evaluate the policy’s language model perplexity, and choose the action with the minimum perplexity averaged over tokens.
@@ 5.4 Results and discussion
Table 3: Task completion rate on three benchmarks. We evaluate a behaviour cloning agent Bc, an imitation learning agent using LLM as the expert policy ActPred, and our proposed method Lfm which imitates desirable behaviour identified by a language feedback model. On held-out evaluation environments, Lfm outperforms other methods on all benchmarks. Furthermore, adaptation to the new environments using the trained language feedback models results in significant additional gains (LfmA).
ALFWorld ScienceWorld Touchdown
Bc 62.6 45.8 57.5
ActPred 56.0 39.0 58.0
Lfm 64.1 47.1 59.7
LfmA 74.6 49.3 62.8
Table 3 shows the performance of the policy behaviour cloned from demonstrations Bc, imitation learned from LLMs using action prediction ActPred, and imitation learned from Lfm. For LFMs, we show zero-shot results (Lfm) as well as after one round of adaptation (LfmA).
@@
Table 3 shows that Lfm improves upon the strong behaviour cloning baseline policy Bc in all benchmarks. Table 5 shows examples of LFM-identified desirable behaviour. This shows that LFMs are an effective means to leverage the knowledge in pretrained LLMs for policy improvement in language-grounded environments, which agree with human-identified desirable behaviour. Appendix D also compares GPT-4 to the open-source Llama 2 70B for training feedback models using human evaluation. We find that GPT-4 consistently outperforms Llama 2, which tends to identify spurious desirable behaviour.
@@
Table 4 shows that LFMs trained during language feedback learning can accurately recognize desirable behaviour in new environments. Table 3 shows that imitating this behaviour obtains significant policy improvement across all benchmarks. This shows that LFMs generalize to new environments, which allows for policy adaptation to new environments despite not having demonstrations nor LLM access.
Table 4: Feedback performance measured by F1 score. We label steps the LLMs consider to be productive to be “positive” actions and other steps negative actions. We measure the F1 score of the positive/negative predictions made by the learned LFM using the LLM predictions as ground truth. We observe no significant performance degradation when using a much more detailed feedback model (LfmD) that also provides explanations behind the feedback, summaries of agent behaviour, and strategy suggestions.
ALFWorld ScienceWorld Touchdown
@@
Table 5: Example of detailed language feedback, which uses more LLM interaction to provide human-interpretable feedback. The prompt and output differences between succinct feedback (Table 2) and detailed feedback are highlighted in red. Note that the second row shows shows batched LLM inference, which is used to train a step-wise feedback model. The step-wise inference of the feedback model is shown in the last row. We leave providing summaries and suggestions with the learned feedback model to future work.
@@
LLM output Yes - Step 28: The player locates the candles by arriving at countertop 1, where a candle is available. - Step 29: The player successfully takes candle 3 from countertop 1, demonstrating the ability to collect… #Summary The player is gathering candles and placing them in a drawer, working towards completing the task of putting two… #Improvement The player can improve their strategy by: - Not closing drawers unnecessarily after checking… - Prioritizing finding and collecting the candles first…
@@ 6 Conclusion
7 Broader Impact
This paper presents work on improving instruction following using Language Feedback Models. Potential beneficial societal consequences of this work include the development of cost-effective computer agents that quickly learn to accurately follow human commands. The method we present in this work learns an language feedback model trained by exploiting world knowledge in LLMs. We show that this technique results in faster and more cost-effective policy improvement than using LLMs as experts. Furthermore, we show that the feedback models we learn generalize to new environments, which results in significant gains for new environments not seen during training via policy adaptation. Finally, we show that language feedback models can be extended to provide detailed critique that include explanations, agent intent summaries, and high-level strategy recommendations. Learning to provide this detailed output results in no noticeable accuracy cost, and can provide interpretable feedback humans can inspect and verify to create more trustworthy policies.
Potential negative societal consequences of this work include hallucinations by LLMs that mislead feedback model training. In this sense, feedback models may learn to encourage actions that do not achieve language goals (e.g. exploring the bathroom during a kitchen cleaning task). Furthermore, they may encourage actions that help achieve goals but are undesirable in other ways (e.g. unsafely climbing over the table to reach the sofa more quickly). In future work, we will explore using techniques in LLM alignment to learn more robust language feedback models, as well as investigate learning from detailed feedback models with human verification to improve the trustworthiness of downstream policies.
Appendix A Verbalization of visual environments

### v6

@@ 4 Related Work
Instruction following in grounded environments has been explored in settings such as navigation [11, 18, 12], game-playing [3, 48], and robotics [8, 37, 9]. However, most prior work model environment observations separately from language instructions by using specialized encoders (e.g. ResNet [19]Bert [16],  Clip [30]), then learn from data how to associate raw observations with language instructions. Instead of solely using raw observations, more recent work verbalize raw observations to describe environments in language [38, 49, 34]. In doing so, observations and instructions can be directly jointly reasoned over using language models to achieve more efficient and generalizable learning through large-scale pretraining. We build on this last direction by verbalizing raw observations into language descriptions to train language policies. However, unlike prior work that train language models to predict next actions, we develop language feedback models that critique verbalized observations and behaviour.
@@
LLMs exhibit reasoning abilities after pretraining on vast quantities of text [10, 43]. A number of recent work on LLMs language agents exploit this reasoning ability. Nakano et al. [28]Yao et al. [46] Deng et al. [15] train instruction following language agents to interact with web browsers to answer questions or interact with web pages. Ahn et al. [2] show that a language agent can be connected with verbalized robots via API interfaces for robotic control. Xie et al. [44] use large visual language models (VLMs) for instruction following in virtual machines, but show that LLMs with verbalization outperform VLMs. While powerful, these prior work are limited in that they require querying an expensive LLM on-line. In contrast, our work examines settings where an LLM is not available on-line. Specially, we use LLMs to collect a small set of off-line data for training LFMs. The small and cost-effective LFMs are then used to identified desirable behaviour for on-line policy improvement without additional interactions with the LLM.
Table 1: Examples of verbalization. We abbreviate long verbalized observations using “…”.
Benchmark Context Action
ALFWorld Task: heat some egg and put it in diningtable. Observation: You arrive at loc 12. On the sinkbasin 1, you see… T-1 Observation: You are in the middle… Action: go to sinkbasin 1 T-2 Observation: … go to microwave 1
ScienceWorld Task: Your task is to find a(n) living thing. First, focus on the thing. Then, move it to the purple box in the bathroom. Observation: You move to the kitchen. This room is called the kitchen. In it, you see: | the agent | a substance called air | a chair. On the chair is… In your inventory, you see: | an orange… T-1 Observation: The door is now open. Action: go to kitchen T-2 Observation… Action: open door to kitchen open door to outside
Touchdown Task: Follow the flow of traffic, with the row of flowers on your left and make a left at the intersection. There will be a white billboard… Observation: behind you, you see: the right lane intersection, a large… T-1 Observation: behind you, slightly… Action: slightly to your left … straight ahead
@@
Recent work enhance language agents by augmenting them with feedback. Ziegler et al. [50]Stiennon et al. [39], and Bai et al. [5] learn reward models from human preference to improve policies via reinforcement learning (RL). Instead of using human feedback, Bai et al. [5] and Lee et al. [24] use LLM feedback to train a separate reward model for RL for textual alignment. Huang et al. [21]Yao et al. [46], and Shinn et al. [35] use LLMs to reason about potential resolutions to failed actions. Yuan et al. [47] use LLMs to generate new prompts and corresponding responses, then use an LLM reward model to identify good prompt-response pairs for self-improvement in text generation alignment. Unlike these approaches, we do not use LLMs during on-line policy improvement. We train an initial small language feedback model from offline LLM data, then use this small feedback model for policy improvement. Additionally, we focus on-line improvement via language feedback for long-horizon, sparse reward, grounded environments instead of text generation alignment. Our procedure for batched, on-line imitation learning is similar to Dagger [31], which we compare to in Appendix F. However, we collect batched expert feedback to identify desirable behaviour instead of corrective actions. Klissarov et al. [22] and Du et al. [17] are recent works that describe learning from feedback approaches complementary to ours. The former learns preference models based on pairwise observations while the latter uses LLMs to suggest exploratory goals during training. Unlike these works, which assume that the underlying goal is the same between training and inference, we consider settings where training and evaluation goals are different. That said, one can expand these approaches to generalize to unseen environments by adapting a preference model during inference [22] and by goal-conditioned subgoal generation during inference [17]. However, unlike LFM, these modifications would then rely on calling LLMs during inference.
@@ 5 Experiments and Analysis
We evaluate using LFM s for policy improvement on three distinct language grounding benchmarks. Formally, the environments from a benchmark are distinct partially-observed Markov Decision Processes that share some (or all) of the environment dynamics but have different instructions, observations, and/or action space.
@@ 5.1 Evaluation benchmarks
Table 1 shows examples of verbalized environments and tasks from each benchmark. Each benchmark provides distinct training and test environments to test generalization. In each environment, the agent takes actions to perform tasks outlined in a language instruction. The task is considered completed if and only if the agent solves the tasks within the preallocated number of steps. We evaluate using task-completion rate over test environments. The statistics from each benchmark is shown in Appendix D Table 6. These three benchmarks share challenges in sparse, delayed reward, partial observability, and compositional generalization to unseen tasks and environments.
ALFWorld is a verbalization of ALFRED [36], a natural language instruction following benchmark set in a 3D simulated kitchen. Here, the agent interacts with objects in kitchens to achieve compositional goals such as cleaning then microwaving potatoes. In ALFWorld [38], raw state information from ALFRED are used to populate language templates that describe observations in language.
ScienceWorld is a textual simulation benchmark for basic science experiments [42]. The agent interacts with objects to conduct experiments specified in natural language, such as determining the boiling temperature of a material. ScienceWorld is uniquely challenging to due the large amount of variations in task types (30), and parametric variations (10-1400) such as the specific substance to be melted. Furthermore, ScienceWorld has a substantially larger action space and longer horizon tasks.
Touchdown is a navigation benchmark where the agent navigates Google Street View images to follow long, compositional instructions [12]. Touchdown requires jointly reasoning over natural images from Google Streetview with occlusion and multi-sentence natural language instructions that describe long-horizon goals. We introduce a new verbalization procedure for Touchdown based on matching noun-phrases and image patches with Clip embeddings to populate egocentric language templates. Behaviour cloning using our verbalization is detailed in Appendix D. Touchdown considers multiple subtasks, in this work we only test the agent’s ability to arrive at the correct location according to the instruction.
@@ 5.2 Methods
Table 2: LLM prompts used to collect desirable behaviour. ActPred uses LLMs to directly generate actions for each step, whereas Lfm uses LLMs to generate batch feedback that identify which taken actions were productive. For brevity, we abbreviate long verbalized observations using “…”. “Before” contains the observation before the first step in the batch.
@@
Prompt Your task is: look at alarmclock under the desklamp. You see: you are in the middle of a room. looking quickly around you, you see a bed 1… what do you decide to do? available actions: examine shelf 1, examine shelf 2, go to bed… You decide to: go to desk 1. You see: you arrive at desk 1. what do you decide to do? available actions: examine desk… You decide to:
LLM Output examine desk 1
@@
Prompt You will be shown a playthrough for solving a task. Task: put two candle in drawer. Before: You open the drawer 6. The drawer 6 is open. In it, you see nothing. Step 27. Your action: close drawer 6. Result: You close the drawer 6… Step 28. Your action: put candle 3 in/on drawer 1. Result: You put the candle 3 in… Is the player on the right track to solve the task? Answer yes or no. If yes, list the helpful steps by the step number in bullet form.
LLM Output Yes - Step 28 - Step 29…
We train BC baseline policies using existing demonstrations. We examine three different techniques for improving the BC policy. Table 2 shows examples of LLM prompts used for each technique.
@@
We learn a small and cost-effective feedback model described in Section 3.2 to identify desirable behaviour for imitation learning. First, we learn a feedback model on the training environments. Second, we use the feedback model to identify desirable behaviour in the training environments for policy improvement via imitation learning. To collect LLM feedback for training LFMs, we collect one rollout for each environment in a benchmark and sample 10k 20-step windows from the rollouts. Crucially, we limit the amount of feedback data collected from the LLM such that the number of output tokens produced by the LLM is identical to ActPred (we use 100k GPT-2 tokens for all benchmarks). For LFM, we collect feedback for as many windows as possible until we exceed 100k output tokens, then use this feedback to train the LFM. For ActPred, we label actions until we exceed 100k output tokens, then combine this labeled set with demonstrations to train the policy. This limitation on LLM interactions answers whether the feedback model is more cost-effective than direct action prediction for imitation learning.
LfmA: adaptation using feedback models.
@@
Table 3: Task completion rates of behaviour cloning Bc, imitation learning (IL) using LLM expert ActPred, and IL using Lfm. On held-out test environments, Lfm outperforms other methods on all benchmarks. ActPred and Lfm are limited to 100k output tokens of GPT-4 interactions. Further adaptation to the new environments using LFM results in significant additional gains (LfmA). Errors are standard deviations across 3 seeds. Previous SOTA are Micheli and Fleuret [27] for ALFWorld, Lin et al. [25] for ScienceWorld, and Schumann and Riezler [33] for Touchdown. Unlike Lin et al. [25], our methods do not use ScienceWorld-specific custom room tracking nor action reranking.
ALF SciWorld TD
Prev SOTA 57.6 45.8 29.3
GPT-4 zeroshot 3.0 1.3 3.2
Bc $62.6\pm 0.4$ $45.8\pm 0.6$ $57.5\pm 0.3$
ActPred $56.0\pm 0.7$ $39.0\pm 0.7$ $58.0\pm 0.4$
Lfm $\textbf{64.1}\pm 0.3$ $\textbf{47.1}\pm 0.5$ $\textbf{59.7}\pm 0.4$
LfmA 1 rnd $74.6\pm 1.1$ $49.3\pm 0.9$ $62.8\pm 1.1$
LfmA 2 rnds $76.5\pm 1.3$ $50.4\pm 1.0$ $63.5\pm 1.2$
@@ 5.3 Experiment details
We use the GPT-4 (2023-03-15) for action prediction and feedback, and finetune 770M Flan-T5 [13] for policy and feedback models. Verbalized observations $v$ contain the most recent 20 steps. We train models for 10k steps with batch 20, learning rate 5e-5, and early stopping over validation demos. For ActPred and LFM, we limit the amount of LLM usage to 100k GPT-2 tokens. Touchdown verbalization uses vit-large-patch14. Appendix H details GPU usage.
@@
We train LFMs using LLM feedback over 20-step windows. We then parse feedback to identify whether the action taken in each step was productive to solving the tasks outlined in the instructions. We subsample the feedback data to obtain an even split of productive and not-productive actions. This data is split into a 80% train/20% validation dataset to train the LFM.
@@
To train policies, we fine-tune language models to minimize token-wise cross-entropy of the ground-truth verbalized action. During inference, we consider a (potentially very large) set of plausible actions given by the environment. For each action, we evaluate the policy’s language model perplexity, and choose the action with the minimum perplexity averaged over tokens.
@@ 5.4 Results and discussion
Table 3 shows the performance of the policy behaviour cloned from demonstrations Bc, imitation learned from LLMs using action prediction ActPred, and imitation learned from Lfm. For LFMs, we show zero-shot (Lfm) as well as adaptation (LfmA) results.
@@
Table 3 shows that Lfm improves upon the strong behaviour cloning baseline policy Bc in all benchmarks. Table 5 shows examples of LFM-identified desirable behaviour. This shows that LFMs are an effective means to leverage the knowledge in pretrained LLMs for policy improvement in language-grounded environments, which agree with human-identified desirable behaviour. Appendix G also compares GPT-4 to the open-source Llama 2 70B for training feedback models using human evaluation. We find that GPT-4 consistently outperforms Llama 2, which tends to identify spurious desirable behaviour.
@@
Table 4 shows that LFMs trained during language feedback learning can accurately recognize desirable behaviour in new environments. Table 3 shows that imitating this behaviour obtains significant policy improvement across all benchmarks. This shows that LFMs generalize to new environments, which allows for policy adaptation to new environments despite not having demonstrations nor LLM access. Appendix I shows additional experiments that demonstrate correlation between LLM/LFM feedback and true state values.
Table 4: Feedback performance of LFM. We measure F1 score of the productive/not-productive predictions made by the learned LFM using the LLM predictions as ground truth. We observe no significant performance degradation when using a much more detailed feedback model (LfmD) that also provides explanations behind the feedback, summaries of agent behaviour, and strategy suggestions.
ALF SciWorld TD
@@
Table 5: Example of detailed human-interpretable feedback. The prompt and output differences between succinct feedback (Table 2) and detailed feedback are highlighted in red. Note that the second row shows shows batched LLM inference, which is used to train a step-wise feedback model. The step-wise inference of the feedback model is shown in the last row. We leave providing summaries and suggestions with the learned feedback model to future work.
@@
LLM output Yes - Step 28: The player locates the candles by arriving at countertop 1, where a candle is available. - Step 29: The player successfully takes candle 3 from countertop 1, demonstrating the ability to collect… #Summary The player is gathering candles and placing them in a drawer, working towards completing the task of… #Improvement The player can improve their strategy by: not closing drawers unnecessarily…
@@ 6 Conclusion
Appendix A Limitations
This work proposes using LLM feedback to improve policies for long-horizon planning in grounded environments. It assumes access to a verbalization module that faithfully describes observations in language. For some practical problems (e.g. rich scenes in operating systems Xie et al. [45]), verbalization has been shown to be a difficult problem, which limits the capability of language feedback models. Moveover, LLMs have been shown to hallucinate, especially in grounded settings that are uncharacteristic in internet pretraining data [42]. Consequently, LLMs may provide inaccurate feedback, which limits policy improvement gains.
Our empirical results demonstrate consistent policy improvement using LFMs on three distinct language grounding benchmarks. These benchmarks cover kitchen interactions (AFLWorld), scientific experiments (ScienceWorld), and real-scene navigation (Touchdown). Future work should investigate the application of LLM feedback learning to real-world robotics, as an important test-bed for LLM ability to provide precise feedback on real-world observations.
While this work significantly reduces the computational burden of using language models as policies in grounded environments by extracting world knowledge from LLMs into small LMs using language feedback, the LM policies we use still require large GPUs to train. Specifically, we note significant performance degradation (e.g. 10% task completion) when changing from 770M Flan-T5 to the smaller 250M variant. Future work should investigate techniques to further reduce model size so that LMs can be tractably used as policies on small devices such as phones.
Appendix B Broader Impacts
Although this work develops LM as policies in grounded environments, the outputs these policies generate are specific (i.e. plans) to the benchmark environments used. Consequently, the generated outputs are not suitable for malicious use (e.g. disinformation, fake profiles). One potential misuse of our proposed method lies in a central party providing a malicious feedback model, which provides negative reinforcement to good behaviour and positive reinforcement to bad behaviour. Learning from such a feedback model can potentially result in malicious downstream policies.
Appendix C Licenses for existing assets
We will release our code and resources under the MIT license. This work uses assets from the Flan-T5 series of models, OpenAI GPT-4, Llama-2, and the three benchmarks. The licenses are as follows:
Appendix D Verbalization of visual environments